OpenStudy (anonymous):

The St. Mark’s Community Barbecue served 250 dinners. A child’s plate cost $3.50 and an adult’s plate cost $7.00. A total of $1347.50 was collected. How many of each type of plate was served?

OpenStudy (gw2011):

Let C = Child's Plate Let A = Adult's Plate C + A = 250 3.5C + 7A = 1347.50 Multiply the first equation by 7 to find what C is and you get: 7C + 7A = 1750 3.5C + 7A = 1347.50 Subtract this equation from the first one and you get: 3.5C = 402.50 C = 115 Substitute this value into the following equation: C + A = 250 115 + A = 250 A = 135 Therefore, 115 Child's Plates and 135 Adult's Plates were served